Consider a system of three equal-mass particles mo
Answer #1

A) position vectors

r1 = (3t^2 +7)i + 0j

r2 = (5t +5)i + 4j

r3 = (4t)i + (3t +5) j

centre of mass position vector

Rcm = (m1r1 + m2r2 + m3r3)/(m1 + m2 +m3) = (r1 +r2 +r3 )/3 = ((3t^2 +7)i + (5t +5)i + 4j +(4t)i + (3t +5) j )/3

=> Rcm = [(3t^2 +9t +12)i + (3t + 9)j ]/3 = (t^2 + 3t +4 )i + (t+3)j m

B) velocity vector

Vcm = dRcm/dt = (2t+3)i + 1j m/s

C) acceleration vector

Acm = dVcm/dt = 2i m/s^2